# Integrating Canonical Neural Units and Multi-Scale Training for Handwritten Text Recognition

### v1

@@ IV-B1 Ablation Experiments
Based on the optimal frame length and single-scale training, Table IV lists the results with/without the 3D attention module and the global-local context information. If we did not adopt the 3D attention module in experiments, a pooling operation was directly employed to reduce the height of the last convolutional feature maps to 1. The 3D attention module can steadily improve the Chinese recognition performances. On the SCUT-HCCDoc data set, the AR is improved from 88.51% to 88.8%. On the SCUT-EPT data set, the AR increases from 75.00% to 76.67%. Although it seems that the English recognition without the 3D attention module has a higher AR (94.31% vs. 94.17%), the network based on the multi-resolutions training still obtains the best AR. As the conclusions demonstrated by many previous researches, there exists an obvious gap with/without using the context information for the sequence task. The results dramatically decrease from 88.8% to 85.36%, 76.67% to 72.28% and 94.17% to 87.0%, respectively. It is interesting to observe that the pure CNN-based models without the 3D attention block and the context information can also achieve 85.76%, 72.53% and 85.93%, respectively.
TABLE IV: The recognition performance with/without different blocks.
@@
In Table V, we compare the results of different training strategies, i.e. the weights are only optimized by the CTC loss or jointly trained by the CTC loss and the CE loss. Although the Chinese decoding results of the auxiliary network are not good, the main recognition network still benefits from the joint training. The best network can achieve 89.34%, 77.15% and 94.41%, respectively. Moreover, we also list the storage and the running time comparisons of different methods. In order to make a fair comparison, all experiments were evaluated on the same machine and the experimental configurations on the same dataset were consistent. The decoding time of the CTC is defined as 1. Generally speaking, the results show that the ED method needs more time consumption and storage.
TABLE V: The comparisons of accuracy rate, storage and relative speed.
@@
IV-B2 Overall Comparison
Finally, Table VI shows an overall comparison of our proposed method and other state-of-the-art methods on different test sets. Without using external data, our proposed network can make a new milestone. Compared with the best networks trained by using external data, the proposed network can also keep comparable results on all datasets. For the listed results on the IAM dataset, the standard partition and no word lexicons were used.
TABLE VI: Performance comparison of our proposed method and other state-of-the-arts methods. For the English handwritten recognition task, all listed results do not use any lexicons.
@@ IV-B2 Overall Comparison
IV-B3 Visualization Analysis
Figure 6: The training loss steadily decreases along with the training epoch.
Figure 7: The character error steadily decreases along with the training epoch.
Fig. 6 and Fig. 7 show the training loss and the character error keep decreasing along with the training epoch, respectively. They both steadily converge into a small interval. In Fig. 8, Fig. 9 and Fig. 10, the attention weights in the 3D block are shown on the original images. It is interesting to observe that even for long-text images and some small symbols, most attention weights (red parts) are still located on the handwriting characters, which demonstrates the proposed attention mechanism is reasonable.
@@ IV-B3 Visualization Analysis
V Conclusion
In this paper, inspired by the segmentation-free methods, we propose a new recognition network by using a novel 3D attention module and global-local context information. Main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are ingeniously organized into a network. The network weights are effectively optimized by the multi-scale training. Experiments on the latest Chinese handwritten text datasets and the English handwritten text dataset show that the proposed method can achieve comparable results with the state-of-the-art methods. Besides, the visualization analysis demonstrates the proposed attention mechanism is reasonable. For future work, we will aim to compress and accelerate the network to reduce decoding time and complete the actual deployment by using more training data and combining the text detection. Furthermore, we will investigate large vision models.
VI Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 62106031. The author would like to thank Professor Jun Du of the University of Science and Technology of China (USTC). When the author was a Ph.D. candidate at USTC, we successfully applied the HMM and the deep neural network to the text recognition task. The supervisor (Prof. Jun Du) always said, can we combine three mainly segmentation-free methods? Although the HMM is not explicitly used in this paper, the question inspired the author to delve into this problem while self-isolated at home.

### v2

@@ 3.2.1 Ablation Experiments
Based on the optimal frame length and single-scale training, Table 4 lists the results with/without the 3D attention module and the global-local context information. If we did not adopt the 3D attention module in experiments, a pooling operation was directly employed to reduce the height of the last convolutional feature maps to 1. The 3D attention module can steadily improve the Chinese recognition performances. On the SCUT-HCCDoc data set, the AR is improved from 88.51% to 88.8%. On the SCUT-EPT data set, the AR increases from 75.00% to 76.67%. Although it seems that the English recognition without the 3D attention module has a higher AR (94.31% vs. 94.17%), the network based on the multi-resolutions training still obtains the best AR. As the conclusions demonstrated by many previous researches, there exists an obvious gap with/without using the context information for the sequence task. The results dramatically decrease from 88.8% to 85.36%, 76.67% to 72.28% and 94.17% to 87.0%, respectively. It is interesting to observe that the pure CNN-based models without the 3D attention block and the context information can also achieve 85.76%, 72.53% and 85.93%, respectively.
Table 4: The recognition performance with/without different blocks.
@@
In Table 5, we compare the results of different training strategies, i.e. the weights are only optimized by the CTC loss or jointly trained by the CTC loss and the CE loss. Although the Chinese decoding results of the auxiliary network are not good, the main recognition network still benefits from the joint training. The best network can achieve 89.34%, 77.15% and 94.41%, respectively. Moreover, we also list the storage and the running time comparisons of different methods. In order to make a fair comparison, all experiments were evaluated on the same machine and the experimental configurations on the same dataset were consistent. The decoding time of the CTC is defined as 1. Generally speaking, the results show that the ED method needs more time consumption and storage.
Table 5: The comparisons of accuracy rate, storage and relative speed.
@@
Table 6: Performance comparison of our proposed method and other state-of-the-arts methods. For the English handwritten recognition task, all listed results do not use any lexicons.
@@
3.2.2 Overall Comparison
Finally, Table 6 shows an overall comparison of our proposed method and other state-of-the-art methods on different test sets. Without using external data, our proposed network can make a new milestone. Compared with the best networks trained by using external data, the proposed network can also keep comparable results on all datasets. For the listed results on the IAM dataset, the standard partition and no word lexicons were used. Please note there exists an unofficial partition (RWTH partition) of the IAM dataset. It is widely used by many researchers, their results are not compared in this paper.
3.2.3 Visualization Analysis
In Fig. 6, Fig. 7 and Fig. 8, the attention weights in the 3D block are shown on the original images. It is interesting to observe that even for long-text images and some small symbols, most attention weights (red parts) are still located on the handwriting characters, which demonstrates the proposed attention mechanism is reasonable.
@@ 3.2.3 Visualization Analysis
4 Conclusion
In this paper, main canonical neural units including attention mechanisms, fully-connected layer, recurrent unit and convolutional layer are ingeniously organized into a network. The network weights are effectively optimized by the multi-scale training. Experimental results show that the proposed method can achieve comparable results with the state-of-the-art methods. Besides, the visualization analysis demonstrates the proposed attention mechanism is reasonable. For future work, we will investigate large vision and language models for document analysis.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 62106031. The author would like to thank Professor Jun Du of the University of Science and Technology of China (USTC). When the author was a Ph.D. candidate at USTC, we successfully applied the HMM and the deep neural network to the text recognition task. The supervisor (Prof. Jun Du) always said, can we combine three mainly segmentation-free methods? Although the HMM is not explicitly used in this paper, the question inspired the author to delve into this problem while self-isolated at home. The author also would like to thank Yan Wang for applying the code of extracting text lines from original SCUT-HCCDoc documents.